An aqueous solution of Nitric acid has an osmotic pressure of 0.80 atm at 27 degree Celsius. The density of solution is 1.03g/mol. Compute for its boiling point at 20 degree Celsius?
1
Expert's answer
2020-09-24T11:22:23-0400

The osmotic pressure "\pi" and the boiling point elevation "\Delta T_b" are the colligative properties of the solutions. The boiling point of the solution can be calculated from:

"\Delta T_b = iK_bm",

where "i" is the van't Hoff factor, "K_b" is the ebullioscopic constant and "m" is the molality of the solution.

On the other hand, the osmotic pressure of the solution can be calculated from:

"\pi = cRTi" ,

where "c" is the molarity, "R" is the gas constant (8.314 J K-1 mol-1) and "T" is the temperature in kelvin (27+273.15 = 300.15 K).

The nitric acid is a strong electrolyte, therefore we can assume that it dissociates completely with the formation of 2 ions:

HNO3 "\rightarrow" H+ + NO3-.

Thus, its van't Hoff factor "i" equals 2.

Now, one can calculate the concentration of the solution of the nitric acid from the osmotic pressure given:

"c = \frac{\pi}{RTi} = \frac{0.8\cdot101325\text{ J/m}^3}{8.314\text{ J/Kmol}\cdot300.15\text{ K}\cdot2} = 16.24 \text{ mol/m}^3" .

The molarity "c" and the molality "m" are linked through the density of the solution and the molar mass of the solute ("M" for HNO3 is 63.01 g/mol):

"m = \left[\frac{d_{sln}}{c} - M_s\right]^{-1} = 0.0158\text { mol/kg}" .

Finally, the boiling point elevation is:

"\Delta T_b = 2\cdot0.512\text{ (°C kg/mol) }\cdot0.0158\text{ (mol/kg)} = 0.016" °C.

Answer: the boiling point elevation is 0.016 °C.